OpenStudy (anonymous):

if using substitution method how do i solve -4x+y=6, -5x+y=21

jimthompson5910 (jim_thompson5910):

Isolate y in the first equation -4x+y=6 -4x+y=6 -4x+y+4x=6+4x y = 4x + 6

jimthompson5910 (jim_thompson5910):

Then plug that into the second equation -5x+y=21 -5x+( y )=21 -5x+( 4x+6 )=21 -5x + 4x + 6 = 21 I'll let you finish up

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome